Prev Next

SW&IT Meeting Minutes 2019-11-18

Attendees: Andrew, Carl, David, Igor, Petr, Stephen, Xianyu, Yang

Lead by: Petr

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20191104
LR20191104 - Bec, Matthieu 80.0
LR20191104 - Holmberg, Carl 40.0
LR20191104 - Colson, Andrew 48.0
LR20191104 - Edmeades, David 48.0
LR20191104 - Sola, Igor 40.0
LR20191104 - Kubanek, Petr 82.7
LR20191104 - Hooper, Stephen 80.0
LR20191104 - Zhang, Yang 70.0
LR20191104 - Zhang, Xianyu 72.0

News from management

Previous Action Items

Open topics

Team reports

PK

  • progress on new GCS - AZCam.cpp (new AzCam interface) able to take images, azcamclient for testing
  • merged GKFP patch from Yang
  • SPIE abstract
  • write some paragraphs for Malaga 2019 paper
  • found minor problems in PCS TimeToLimit - patch for Sunday engineering (unwrapping from position commaned with MCSPU azslewtohold,..)
  • IT7917: PMC issues, searching for cause
  • PCSGUI: PR to show rotators position
  • patched PCS during engineering

XZ

Last week:
  • SPIE paper abstracts
  • TNO ASM meeting
  • IDL license issue
  • AGW2 support
This week:
  • Setup the mountain anallyze VM
  • 4D trigger test
  • IDL license test and support for license server switch
  • Reponse to SOUL FLAO-like operability
  • AO telemetry document
  • SOUL lab setup support?

DE

  • Return KVM test unit
    • Worked well, but I'd rather return it now and get the most appropriate number of ports when we need them
  • VMWare support contract with CDW now due to purchasing restrictions
  • Al VM / F30, all functional from my end, waiting for feedback from him
    • Friday reported all good
  • Vacation
  • Shall we buy test copies of 14TB disks to trial in the Synology before committing to that course of action?
  • Storage VAR came back with a pretty amazing deal on Isilon storage--somehow $70k disappeared from the bottom line
    • ~350TiB usable capacity in 6U
    • Adequate for all current needs
    • Can handle all of the various requests we've been getting recently
  • Investigate Cisco training/certification
  • UITS changed the mailing list for the Solarwinds alert emails

YZ

  • MODS:
    • The MODS GUI problem (IT#7855). Currently having trouble debugging it (seems difficult to create a local simulation).
    • IT#7927 (MODS image corrupted during transfer). Currently working with Olga/Al/LBT archive/OSU to create a solution.
    • Reading some of the previous MODS handover documents. Preparing for further discussion in near future.
  • LBC:
    • Continue to study and work on the fix for the LBC non-sidereal guiding issue, IT#7847 (have not spent too much time on this).
    • Reviewing the LBT OB GUI Java code, potentially cleaning it and then put it in Github.
  • AO-Cravat:
    • Developed a Python script that creates OSCO sequence data based on the Sciarc Mongodb.
    • Continue to develop the web interface to match the new OSCO sequence data format.
  • Scheduled with Olga a time to work through the remote room new workstation on LBC and MODS operations.

AC

  • Finished ARGOS desktop navigation menu
  • ARGOS/TaN testing
  • OAC config file - need to update git version
  • PMS telemetry
  • ARGOS transition

MB

  • OC, WPC, Management
  • DMS zabbix adapter WIP
  • LBTI alarm scheme round up, next actions
  • quick eval CentOS -stream, looking good but will need 3rd party repos for now (EPEL?)
  • look into interesting MODS striping issues, due to file transfer ??
next
  • GMT TMS review
  • Power restart checkout list
  • PEPSI servers, meet and wrap up - orders must go now
  • iLocater snmp get some closure

IS

Last week:
  • miniQ
    • working on the report part, and routing and history navigation.
  • PIT
    • work on ways to improve and automatize proposals after period closes.
    • work on implementing a resubmit option to the PIT so there are no or to limit duplicates
  • LUCI
    • looking at observer compunter GEIRS launch script issues and RTD fits file caching.
This week
  • PIT
    • allocated some time to this.
  • LUCI
    • any issues related to the 64 bit migration

SH

  • installed new vCenter, Louise on VMWare licensing, all mountain VMs off of Synology except for the mountainapp1, reworked "dms" to be "dms", and "archive" to be "archive".
  • moved telescope home on obs1, and LSS/syslog off of Synology
  • iLocater network investigation
  • moved night report tool to write to /lbt/data/logs directory instead of /home/telescope/sallanson
  • AGW2 network in clean room
  • Machines for SOUL work in Tucson

next:

  • RPM packages for CentOS 8 Stream
  • install robs and robs32 mountain, GFS2
  • 507-2 forensics
  • install ovirt on analysis(?), vm1 mountain old

CH

Last week
  • ALH: Prototyped a Google Mail API-based script to retrieve LBTI chiller PLC alert emails. Met with S Ertel and M Montoya to develop requirements.
  • Supportscripts: Began testing RB_Science.py environment to diagnose the DS9 crashes John Hill is seeing when two scripts are communicating on XPA at once.
  • TMS: Added a test to directly test up/down state of TMS host.

This week
  • ALH: Clarify LBTI alerting reqs and sketch out an implementation.
  • ARGOS: Begin learning the system.
  • Supportscripts: Diagnose DS9 crashes during multiple script runs.
  • TMS: Hoping for opportunity to test status script/ALH integration mid week.
Topic revision: r13 - 18 Nov 2019, AndrewColson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback